# Breaking the degeneracy among regular black holes with gravitational lensing

### 21 Mar 2026 v1

@@ 5.1 Subleading observables
5.2 Shadows with spherical accretion flows
We investigate the shadows of regular black holes immersed in both static and infalling spherical accretion flows. Restricted to the equatorial plane ($\theta=\pi/2$), the observed specific intensity is determined by integrating the emissivity along photon trajectories, which accounts for gravitational redshift Jaroszynski:1997bw
@@ 5.2 Shadows with spherical accretion flows
5.3 Probing Regular Black Holes with ngEHT
The high order signatures $\{s,r_{\rm mag},\Delta T^{1}_{n,m},\lambda\}$ established in this work provide a concrete benchmark for the ngEHT. While $b_{m}$ exhibits “macroscopic universality,” these sub-leading corrections are sensitive to the fine grained curvature of non-singular cores. With $\sim 15\,\mu\mathrm{as}$ resolution and $1000:1$ dynamic range Johnson:2023ynn , the ngEHT can resolve $n=1,2$ lensing sub-rings, allowing a direct measurement of the Lyapunov exponent $\lambda$ to distinguish between the unstable orbits of Culetu and the relatively stabilized Hayward geometry. Furthermore, observations at $345GHz$ will exploit a more transparent plasma regime Johnson:2023ynn , bypassing the optically thick layers that obscure near horizon physics. This transparency, combined with high cadence monitoring, enables the detection of the predicted brightness hierarchy inversion in “black hole movies.” Since the Culetu model’s unique lensing velocity alignment results in less severe Doppler suppression, its relative intensity provides a robust channel to break geometric degeneracies and distinguish between specific regular black hole formulations.
6 Conclusion
@@ Appendix A Weak field deflection angle and Einstein ring
In this section, we briefly present the derivation of weak deflection angle $\alpha(b)$ and subsequent determination of Einstein radius $\theta_{E}$ for black hole metrics considered in this work. Following the methodology of Ref. Keeton:2005jd , we perform a post-Newtonian expansion in the weak field limit ($x_{0}\gg 1$) to express the deflection angle as
$\begin{aligned} \alpha(x_{0})=&\frac{4}{x_{0}}+\frac{15\pi/4-4+A_{2}}{x_{0}^{2}}+\frac{122/3-15\pi/2+A_{3}}{x_{0}^{3}}\\ &+\frac{-130+3465\pi/64+A_{4}}{x_{0}^{4}}+\mathcal{O}(x_{0}^{-5}),\end{aligned}$ (29)
@@
$\begin{aligned} \alpha(b)=&\frac{4M}{b}+\left(\frac{15\pi}{4}+B_{2}\right)\frac{M^{2}}{b^{2}}+\left(\frac{128}{3}+B_{3}\right)\frac{M^{3}}{b^{3}}\\ &+\left(\frac{3465\pi}{64}+B_{4}\right)\frac{M^{4}}{b^{4}}+\mathcal{O}(b^{-5}),\end{aligned}$ (30)
@@
$\begin{aligned} \theta_{E4}=&\frac{1}{16\delta^{9/2}}\Big[-6\delta^{2}\gamma^{2}-6\delta^{5/2}\gamma^{2}+12\delta^{3}\gamma^{2}+9\gamma^{3}\\ &-25\delta^{1/2}\gamma^{3}-24\delta\gamma^{3}+48\delta^{3/2}\gamma^{3}+8\delta^{3}\mu\\ &-8\delta^{7/2}\mu+12\delta\gamma\mu+4\delta^{3/2}\gamma\mu-32\delta\gamma\mu+8\delta^{5/2}\nu\Big].\end{aligned}$ (37)
@@ Appendix B The Bozza’s procedure
In this appendix we briefly review Bozza’s analytical formalism for strong gravitational lensing Bozza:2002zj . A photon orbiting a black hole has an innermost unstable circular orbit known as the photon sphere, with radius $x_{m}$. The impact parameter $b$ is conserved along the trajectory and relates to the closest approach $x_{0}$ as $b=\sqrt{C_{0}/A_{0}}$, where the subscript “0” denotes evaluation at $x_{0}$. The critical impact parameter corresponds to the photon sphere and is given by $b_{m}=\sqrt{C_{m}/A_{m}}$, where “mm” indicates evaluation at $x_{m}$. Introducing the variable $y=A(x)$ and $z=(y-y_{0})/(1-y_{0})$, the deflection integral can be rewritten as
@@
$\begin{aligned} R(z,x_{0})&=\frac{2\sqrt{By}}{CA^{\prime}}(1-y_{0})\sqrt{C_{0}},\\ f(z,x_{0})&=\frac{1}{\sqrt{y_{0}-[(1-y_{0})z+y_{0}]C_{0}/C}}\end{aligned}$ (39)
@@
$\begin{aligned} \tilde{\beta}=&\frac{(1-y_{0})^{2}}{2C_{0}^{2}{A^{\prime}_{0}}^{3}}\Big[2C_{0}C^{\prime}_{0}{A^{\prime}_{0}}^{2}+(C_{0}C^{\prime\prime}_{0}-2{C^{\prime}_{0}}^{2})y_{0}A^{\prime}_{0}-\\ &\quad C_{0}C^{\prime}_{0}y_{0}A^{\prime\prime}_{0}\Big].\end{aligned}$ (42)
@@
$\begin{aligned} \tilde{I}_{D}(x_{0})&=\int_{0}^{1}R(0,x_{m})f_{0}(z,x_{0})dz,\\ \tilde{I}_{R}(x_{0})&=\int_{0}^{1}\big(R(z,x_{m})f(z,x_{m})-R(0,x_{m})f_{0}(z,x_{m})\big)dz.\end{aligned}$ (44)
@@
$\begin{aligned} \bar{a}&=\frac{R(0,x_{m})}{2\sqrt{\beta_{m}}},\\ \bar{b}&=-\pi+\bar{a}\ln\frac{2\beta_{m}}{y_{m}}+b_{R},\end{aligned}$ (47)

### 26 Mar 2026 v2

@@ V.1 Subleading observables
V.2 Shadows with spherical accretion flows
We investigate the shadows of regular black holes immersed in both static and infalling spherical accretion flows. Restricted to the equatorial plane ($\theta=\pi/2$), the observed specific intensity is determined by integrating the emissivity along photon trajectories, which accounts for gravitational redshift [92]
@@ V.2 Shadows with spherical accretion flows
V.3 Probing regular black holes with ngEHT
The high order signatures $\{s,r_{\rm mag},\Delta T^{1}_{n,m},\lambda\}$ established in this work provide a concrete benchmark for the ngEHT. While $b_{m}$ exhibits “macroscopic universality,” these sub-leading corrections are sensitive to the fine grained curvature of non-singular cores. With $\sim 15\,\mu\mathrm{as}$ resolution and $1000:1$ dynamic range [85], the ngEHT can resolve $n=1,2$ lensing sub-rings, allowing a direct measurement of the Lyapunov exponent $\lambda$ to distinguish between the unstable orbits of Culetu and the relatively stabilized Hayward geometry. Furthermore, observations at $345GHz$ will exploit a more transparent plasma regime [85], bypassing the optically thick layers that obscure near horizon physics. This transparency, combined with high cadence monitoring, enables the detection of the predicted brightness hierarchy inversion in “black hole movies.” Since the Culetu model’s unique lensing velocity alignment results in less severe Doppler suppression, its relative intensity provides a robust channel to break geometric degeneracies and distinguish between specific regular black hole formulations.
VI Conclusion
@@ Appendix A WEAK FIELD DEFLECTION ANGLE AND EINSTEIN RING
In this section, we briefly present the derivation of weak deflection angle $\alpha(b)$ and subsequent determination of Einstein radius $\theta_{E}$ for black hole metrics considered in this work. Following the methodology of Ref. [27], we perform a post-Newtonian expansion in the weak field limit ($x_{0}\gg 1$) to express the deflection angle as
$\alpha(x_{0})=$ $\frac{4}{x_{0}}+\frac{15\pi/4-4+A_{2}}{x_{0}^{2}}+\frac{122/3-15\pi/2+A_{3}}{x_{0}^{3}}$
$+\frac{-130+3465\pi/64+A_{4}}{x_{0}^{4}}+\mathcal{O}(x_{0}^{-5}),$ (29)
@@
$\alpha(b)=$ $\frac{4M}{b}+\left(\frac{15\pi}{4}+B_{2}\right)\frac{M^{2}}{b^{2}}+\left(\frac{128}{3}+B_{3}\right)\frac{M^{3}}{b^{3}}$
$+\left(\frac{3465\pi}{64}+B_{4}\right)\frac{M^{4}}{b^{4}}+\mathcal{O}(b^{-5}),$ (30)
@@
$\theta_{E4}=$ $\frac{1}{16\delta^{9/2}}\Big[-6\delta^{2}\gamma^{2}-6\delta^{5/2}\gamma^{2}+12\delta^{3}\gamma^{2}+9\gamma^{3}$
$-25\delta^{1/2}\gamma^{3}-24\delta\gamma^{3}+48\delta^{3/2}\gamma^{3}+8\delta^{3}\mu-8\delta^{7/2}\mu$
$+12\delta\gamma\mu+4\delta^{3/2}\gamma\mu-32\delta\gamma\mu+8\delta^{5/2}\nu\Big].$ (37)
@@ Appendix B THE BOZZA’S PROCEDURE
In this appendix we briefly review Bozza’s analytical formalism for strong gravitational lensing [32]. A photon orbiting a black hole has an innermost unstable circular orbit known as the photon sphere, with radius $x_{m}$. The impact parameter $b$ is conserved along the trajectory and relates to the closest approach $x_{0}$ as $b=\sqrt{C_{0}/A_{0}}$, where the subscript “0” denotes evaluation at $x_{0}$. The critical impact parameter corresponds to the photon sphere and is given by $b_{m}=\sqrt{C_{m}/A_{m}}$, where “mm” indicates evaluation at $x_{m}$. Introducing the variable $y=A(x)$ and $z=(y-y_{0})/(1-y_{0})$, the deflection integral can be rewritten as
@@
$R(z,x_{0})=\frac{2\sqrt{By}}{CA^{\prime}}(1-y_{0})\sqrt{C_{0}},$ (39)
$f(z,x_{0})=\frac{1}{\sqrt{y_{0}-[(1-y_{0})z+y_{0}]C_{0}/C}}$ (40)
@@
$\tilde{\beta}$= $\frac{(1-y_{0})^{2}}{2C_{0}^{2}{A^{\prime}_{0}}^{3}}\Big[2C_{0}C^{\prime}_{0}{A^{\prime}_{0}}^{2}+(C_{0}C^{\prime\prime}_{0}-2{C^{\prime}_{0}}^{2})y_{0}A^{\prime}_{0}$-
$\quad C_{0}C^{\prime}_{0}y_{0}A^{\prime\prime}_{0}\Big].$ (43)
@@
$\tilde{I}_{D}(x_{0})=\int_{0}^{1}R(0,x_{m})f_{0}(z,x_{0})dz,$ (45)
$\tilde{I}_{R}(x_{0})=\int_{0}^{1}\big(R(z,x_{m})f(z,x_{m})-R(0,x_{m})f_{0}(z,x_{m})\big)dz.$ (46)
@@
$\bar{a}=\frac{R(0,x_{m})}{2\sqrt{\beta_{m}}},$ (49)
$\bar{b}=-\pi+\bar{a}\ln\frac{2\beta_{m}}{y_{m}}+b_{R},$ (50)